# Elliptic curves and Fourier coefficients of meromorphic modular forms

### v1

@@ Abstract.
We discuss several congruences satisfied by the coefficients of meromorphic modular forms, or equivalently, $p$-adic behaviors of meromorphic modular forms under the $U_{p}$ operator, that are summarized from numerical experiments, connecting meromorphic modular forms to symmetric powers of elliptic curves. We also provide heuristic explanations for these congruences as well as prove some of them using hypergeometric functions and the Borcherds–Shimura lift.
@@
meromorphic modular forms, symmetric powers of elliptic curves, magnetic modular forms, supercongruences
@@ 1. Introduction
Coefficients of modular forms have always been one of the central topics in the field of modular forms. The vast majority of work in this field is restricted to the study of holomorphic or weakly holomorphic modular forms. In this paper, we instead turn to meromorphic modular forms. The main objects of interest are meromorphic modular forms of level $1$ with precisely one pole at a non-cuspidal point. Specifically, we will focus on modular forms of the form
(1) $A_{1}\cdot\frac{g}{j-c}+A_{2}\cdot\frac{g}{(j-c)^{2}}+\cdots+A_{r}\cdot\frac{g}{(j-c)^{r}},$
where $c\in\mathbb{Q}$, each $A_{i}\in\mathbb{Q}$, and $g\in\mathbb{Z}[[q]]$ is a holomorphic modular form of level $1$ such that $\frac{g}{j-c}$ has a pole. We will discuss the (conjectural) $p$-adic properties possessed by these meromorphic modular forms as well as the heuristic reasons for those properties. In particular, we hope to convince readers that these meromorphic modular forms of weight $k$ behave similarly to symmetric (k2)(k-2)-nd powers of elliptic curves. Indeed, the paper by Brown–Fonseca [BF25] already evidences such a connection.
In this introduction, we will motivate the subject through several phenomena satisfied by the modular form $\frac{E_{4}}{j-c}$ for some $c\in\mathbb{Q}$, some provable and some conjectural. We start with the cases when $c\in\{0,1728\}$, i.e., $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$. Here we make a definition that a (general) modular form $\sum_{n=1}^{\infty}a_{n}q^{n}\in\mathbb{Z}[[q]]$ is called magnetic if $n{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$, and specifically $r$-magnetic for some $r\in\mathbb{Z}_{\geq 0}$ if $n^{r}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$ for all $n\in\mathbb{Z}^{+}$. We have the following result by Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorem 1].
@@
The first numerically observed example of a magnetic modular form appeared in Broadhurst–Zudilin [BZ19, Conjecture 4] with motivation coming from the Hall effect in electromagnetism. The magnetic property of that example, together with that of $\frac{64E_{4}}{j}$, was proved in Li–Neururer [LN19]. Their proof uses the Borcherds–Shimura lift [Bor98]: one identifies the half-integral weight preimage, proves certain divisibility properties of the coefficients of the preimage, and deduces the magnetic property of the image. This method was later generalized by Paşol–Zudilin [PZ22] to show the magnetic property of other meromorphic modular forms, including $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, all of which have poles at CM points. For a detailed discussion of magnetic modular forms, we also refer to the paper by Bönisch–Duhr–Maggio [BDM24] and the references therein.
In fact, by modifying the method of using the Borcherds–Shimura lift, one can show that both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ also satisfy the following supercongruences  [^1] Here the word “supercongruence” refers to the type of congruence with modulus $p^{rl}$ for some $r>1$., which almost imply their $1$-magnetic property and in particular confirm that they are not $2$-magnetic. In Section 4, we will illustrate that both the magnetic property and the supercongruences occur in general for meromorphic modular forms with CM poles.
@@
It is known that $|a_{p}(C)|\leq 2\sqrt{p}$ by the Hasse–Weil bound. A good prime $p$ of $C$ is called an ordinary (resp. supersingular) prime of $C$ if $p\nmid a_{p}(C)$ (resp. $p{\hskip 1.0pt}|{\hskip 1.0pt}a_{p}(C)$). For $p\geq 5$, this is equivalent to the condition that $a_{p}(C)\neq 0$ (resp. $a_{p}(C)=0$).
@@
Here the appearance of $a_{p}(C)^{2}$ already hints at the link between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$, since the trace of Frobenius at $p$ of $\mathrm{Sym}^{2}C$ is $a_{p}(C)^{2}-p$, which is congruent to $a_{p}(C)^{2}$ modulo $p$. When $p$ is supersingular, Theorem 1.3 a priori only implies that $a_{p}(\frac{E_{4}}{j-j(C)})\equiv 0\pmod{p}$, while in fact numerical observations suggest the following stronger congruence.
@@
In general, the connection between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$ can be visualized via the following longer congruence, where one considers suitable linear combinations of multiple coefficients of $\frac{E_{4}}{j-j(C)}$.
@@
$a_{np}-A_{p}{\hskip 1.0pt}a_{n}+\chi(p)p^{k-1}a_{n/p}\;\equiv\;0\pmod{p^{(k-1)(v_{p}(n)+1)}}$.
In general, it also refers to congruences of the form
@@
for some suitable power of $p$. Atkin and Swinnerton-Dyer [ASD71] first came across this type of congruences when studying the coefficients of some noncongruence modular form, whose result was later generalized by Scholl [Sch85]. In Kazalicki–Scholl [KS16], they also demonstrated certain ASD congruences for weakly holomorphic modular forms, and as an example, showed that $F=E_{4}^{6}/\Delta-1464E_{4}^{3}$ satisfies that
@@
where $\tau$ is the Ramanujan tau function. It has also come to the author’s attention that Michael Allen, Ling Long, and Hasan Saad are approaching the ASD congruences for meromorphic modular forms on finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by extending the methods of Scholl and Kazalicki–Scholl, which may lead to proofs of some of the conjectures described in this paper. For a comprehensive overview of the ASD congruences, we refer to the paper by Li–Long [LL14] and the references therein.
We will end the list of phenomena with an explicit connection between the $p$-th coefficient of $\frac{E_{4}}{j-c}$ and the truncated hypergeometric sum associated to the datum $\big((\frac{1}{2},\frac{1}{6},\frac{5}{6}),(1,1)\big)$
@@
As discussed earlier, the cases when $c\in\{0,1728\}$ behave differently from the rest. This is because both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ have a double pole while $\frac{E_{4}}{j-c}$ for $c\notin\{0,1728\}$ only has a simple pole. In general, the cases when $c\in\{0,1728\}$ usually involve more technicalities. We will thus assume for simplicity that $c\notin\{0,1728\}$ throughout most part of the paper, and only briefly discuss those two excluded cases in Section 7.1. When investigating the $p$-adic properties, we will also assume that $v_{p}(c)=0=v_{p}(c-1728)$ throughout the paper.
The reason that all the phenomena have rather neat formulations in weight $4$ is that there is no cusp form. In general, the coefficients of linear combinations of $\big\{\frac{g}{(j-j(C))^{r}}\big\}_{r}$ are an interaction of the elliptic curve $C/\mathbb{Q}$ and the cusp forms of weight $k$, whose formulation is more complicated. For ease of exposition, we will mostly focus on the cases when $k\in\{4,6,8,10,14\}$ so that there is no influence from cusp forms but that there is still a (canonical) choice for $g$, i.e., the Eisenstein series $E_{k}$ of weight $k$. In Section 7.2, we will describe the way of generalizing the whole setting to other weights.
It should also be noted that while this paper heavily adopts the word “conjectures”, they should be more accurately named as numerical observations, as they are summarized from numerical experiments with these meromorphic modular forms.

### v2

@@ Abstract.
We discuss several congruences satisfied by the coefficients of meromorphic modular forms, or equivalently, the $p$-adic behaviors of meromorphic modular forms under the $U_{p}$ operator, that are summarized from numerical experiments. In the generic case, we observe the connection to symmetric powers of elliptic curves, while in the CM case, we furthermore observe the connection to the $p$-adic analogue of the Chowla–Selberg periods. Along with the discussions, we will provide some heuristic explanations for these congruences as well as prove some of them using hypergeometric functions and the Borcherds–Shimura lift.
@@
meromorphic modular forms, symmetric powers of elliptic curves, magnetic modular forms, supercongruences, Chowla–Selberg periods.
@@ 1. Introduction
Coefficients of modular forms have always been one of the central topics in the field of modular forms. The vast majority of work in this field is restricted to the study of holomorphic or weakly holomorphic modular forms. In this paper, we instead turn to meromorphic modular forms. The main objects of interest are meromorphic modular forms of level $1$ with precisely one pole at a non-cuspidal point. We will discuss the conjectural [^1] It should be noted that while this paper heavily adopts the word “conjectures”, they should be more accurately named as numerical observations, as they are summarized from numerical experiments with these meromorphic modular forms. $p$-adic properties possessed by these meromorphic modular forms as well as the heuristic reasons for those properties. In particular, it turns out that these meromorphic modular forms of weight $k$ behave similarly to symmetric (k2)(k-2)-nd powers of elliptic curves. Indeed, the paper by Brown–Fonseca [BF25] already evidences such a connection. We will also discuss the CM case in detail, namely when the non-cuspidal pole is at a CM point, which may be the most interesting part of this paper and deserves thorough study in its own right. In this case, we will construct a family of meromorphic modular forms whose conjectural $p$-adic behaviors can be characterized via the $p$-adic analogue of the Chowla–Selberg periods, which is then related to special values of $p$-adic $L$-functions via Ferrero–Greenberg [FG79].
For ease of exposition, throughout the first half of this paper, we will mostly focus on the cases of weight $k\in\{4,6,8,10,14\}$, and specifically, on modular forms of the form
$A_{1}\cdot\frac{E_{k}}{j-c}+A_{2}\cdot\frac{E_{k}}{(j-c)^{2}}+\cdots+A_{r}\cdot\frac{E_{k}}{(j-c)^{r}},$
where $c\in\mathbb{Q}$, each $A_{i}\in\mathbb{Q}$, and $E_{k}$ is the Eisenstein series of weight $k\in\{4,6,8,10,14\}$. For general weights, we will first remove the influence of cusp forms using relations, or equivalently, certain linear combinations of the Hecke operators, and then investigate the $p$-adic properties of the resulting modular forms. We will discuss this generalization and prove the main results for all weights in the second half of this paper.
1.1. The case of weight 44
We first motivate the subject through several phenomena satisfied by the modular form
$\frac{E_{4}}{j-c}$
for some $c\in\mathbb{Q}$. We start with the cases when $c\in\{0,1728\}$, i.e., $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$. Here we make a definition that a (general) modular form $\sum_{n=1}^{\infty}a_{n}q^{n}\in\mathbb{Z}[[q]]$ is called magnetic if $n{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$ for all $n\in\mathbb{Z}^{+}$, and specifically $r$-magnetic for some $r\in\mathbb{Z}_{\geq 0}$ if $n^{r}{\hskip 1.0pt}|{\hskip 1.0pt}a_{n}$ for all $n\in\mathbb{Z}^{+}$. We have the following result by Li–Neururer [LN19, Theorem 1.5] and Paşol–Zudilin [PZ22, Theorem 1].
@@ 1.1. The case of weight 44
The first numerically observed example of a magnetic modular form appeared in Broadhurst–Zudilin [BZ19, Conjecture 4] with motivation coming from the Hall effect in electromagnetism. The magnetic property of that example, together with that of $\frac{64E_{4}}{j}$, was proved in Li–Neururer [LN19]. Their proof uses the Borcherds–Shimura lift: one identifies the half-integral weight preimage, proves certain divisibility properties for the preimage, and deduces the magnetic property of the image. This method was later generalized by Paşol–Zudilin [PZ22] to show the magnetic property of other meromorphic modular forms, including $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, all of which have poles at CM points. For a detailed discussion of magnetic modular forms, we also refer to the paper by Bönisch–Duhr–Maggio [BDM24] and the references therein.
By modifying the method of using the Borcherds–Shimura lift, one can also show that both $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$ satisfy the following supercongruences  [^2] Here the word “supercongruences” refers to the type of congruences with modulus $p^{rl}$ for some $r>1$., which almost imply their $1$-magnetic property and in particular confirm that they are not $2$-magnetic. In general, both the magnetic property and the supercongruences occur for meromorphic modular forms with CM poles, as we will see in Section 4.
@@
A good prime $p$ of $C$ is called an ordinary (resp. supersingular) prime of $C$ if $p\nmid a_{p}(C)$ (resp. $p{\hskip 1.0pt}|{\hskip 1.0pt}a_{p}(C)$). For $p\geq 5$, this is equivalent to the condition that $a_{p}(C)\neq 0$ (resp. $a_{p}(C)=0$).
@@
Here the appearance of $a_{p}(C)^{2}$ already hints at the link between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$, since the trace of Frobenius at $p$ on $\mathrm{Sym}^{2}C$ is $a_{p}(C)^{2}-p$, which is congruent to $a_{p}(C)^{2}$ modulo $p$. When $p$ is supersingular, Theorem 1.3 a priori only implies that $a_{p}(\frac{E_{4}}{j-j(C)})\equiv 0\pmod{p}$, while in fact numerical observations suggest the following stronger congruence.
@@
In general, the connection between $\frac{E_{4}}{j-j(C)}$ and $\mathrm{Sym}^{2}C$ can be visualized via the following longer congruence, which involves suitable linear combinations of several coefficients of $\frac{E_{4}}{j-j(C)}$.
@@
$a_{np}-A_{p}{\hskip 1.0pt}a_{n}+\chi(p)p^{k-1}a_{n/p}\;\equiv\;0\pmod{p^{(k-1)(v_{p}(n)+1)}}$
for some sequence $(a_{n})$. In general, it also refers to congruences of the form
@@
for some suitable power of $p$. Atkin and Swinnerton-Dyer [ASD71] first came across this type of congruences when studying the coefficients of modular forms on non-congruence subgroups, whose result was later generalized by Scholl [Sch85]. In [KS16], Kazalicki and Scholl demonstrated certain ASD congruences for weakly holomorphic modular forms, and as an example, showed that $F=E_{4}^{6}/\Delta-1464E_{4}^{3}$ satisfies that
@@
where $\tau$ is the Ramanujan tau function. More recently, Allen, Long, and Saad [ALS25] established certain ASD congruences for meromorphic modular forms on finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$. For an overview of the ASD congruences, we refer to the paper by Li–Long [LL14].
We end the list of phenomena with an explicit connection between the $p$-th coefficient of $\frac{E_{4}}{j-c}$ and the truncated hypergeometric sum associated to the datum $\big((\frac{1}{2},\frac{1}{6},\frac{5}{6}),(1,1)\big)$
@@
1.2. Main results
We now discuss the main results of this paper. We will fix the weight $k\in\{4,6,8,10,14\}$ for ease of exposition, though the results are proved for all weights.
The first result is a generalization of Theorem 1.3 on meromorphic modular forms with a simple pole.
Theorem 1.7.
Let $k\in\{4,6,8,10,14\}$, $C/\mathbb{Q}$ be an elliptic curve with $j(C)\notin\{0,1728\}$, and $p\geq 5$ be a good prime of $C$ with $v_{p}(j(C))=0=v_{p}(j(C)-1728)$. Then, for all $n\in\mathbb{Z}^{+}$,
$a_{np}\bigg(\frac{E_{k}}{j-j(C)}\bigg)\;\equiv\;a_{p}(C)^{k-2}a_{n}\bigg(\frac{E_{k}}{j-j(C)}\bigg)\pmod{p}$.
Remark 1.8.
This theorem is conjecturally true for $p=2,3$. In the case of a general weight $k$, conjecturally one can sometimes also remove the assumption $v_{p}(j(C))=0$ or $v_{p}(j(C)-1728)=0$ depending on $k\pmod{12}$ (see Remark 5.2).
Remark 1.9.
We will prove a suitable generalization of this theorem that holds for all weights and for all elliptic curves over number fields (Theorem 7.1). The key ingredient of proof is the hypergeometric connection, i.e., Theorem 1.6.
The second result is on the supercongruences and the magnetic property of certain meromorphic modular forms of weight $k$ with a CM pole of order $k/2$. In order to describe the result, we first introduce some notations.
Recall that for a (nearly holomorphic) modular form $f(\tau)$ of weight $k_{f}$, we may view $f$ as a function in $\tau$ and $\overline{\tau}$ and define the non-holomorphic derivative as
It is easy to check that $\partial_{\tau}f$ is modular of weight $k_{f}+2$. Thus, it is possible to apply the operator $\partial_{\tau}$ recursively to a modular form.
Let $D<0$ be a discriminant with class number $1$, $\alpha_{D}\in\mathcal{H}$ be a CM point of discriminant $D$, $\mathcal{O}_{D}$ be the imaginary quadratic order of discriminant $D$, and $w_{D}:=|\mathcal{O}_{D}^{\times}|$. For $k\in\{4,6,8,10,14\}$, let  [^3] This two-variable modular form $G_{k}(z,\tau)$ occurred in many places throughout the literature. We refer to [AKN97, p. 98] for a detailed discussion and the references therein. It also plays a key role in Zagier’s duality [Zag02, DJ08a].
$G_{k}(z,\tau)\;:=\;\frac{E_{k}(z)\cdot\frac{E_{14-k}}{\Delta}(\tau)}{j(z)-j(\tau)}$ (1)
and
$\mathcal{G}_{k,D}(z)\;:=\;\frac{2}{w_{D}}\cdot(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\alpha_{D}).$ (2)
Here $G_{k}(z,\tau)$ is modular of weight $k$ and 2k2-k with respect to $z$ and $\tau$ respectively, so $(\partial^{\frac{k-2}{2}}_{\tau}G_{k})(z,\tau)$ is of weight 0 with respect to $\tau$. Hence, $\mathcal{G}_{k,D}(z)$ is independent of the choice of $\alpha_{D}$.
Theorem 1.10.
Let $k\in\{4,6,8,10,14\}$. Let $D<0$ be a discriminant of class number $1$ and write $D=A^{2}D_{0}$ for some $A\in\mathbb{Z}$ and some fundamental discriminant $D_{0}<0$. Let
[…]
Then, $\widetilde{\mathcal{G}}_{k,D}$ satisfies the following.
  1. (1)
    $\widetilde{\mathcal{G}}_{k,D}$ has integer Fourier coefficients and is $\frac{k-2}{2}$-magnetic, i.e., for all $n\in\mathbb{Z}^{+}$,
  2. (2)
    Let $p$ be a prime with $p\nmid A$. Then, for all $n,l\in\mathbb{Z}^{+}$,
    $a_{np^{l}}(\widetilde{\mathcal{G}}_{k,D})\;\equiv\;\big(\big(\tfrac{D}{p}\big)p\big)^{\frac{k-2}{2}}a_{np^{l-1}}(\widetilde{\mathcal{G}}_{k,D})\pmod{p^{(k-1)l}}$,
    where $(\frac{\cdot}{p})$ denotes the Legendre symbol.
Remark 1.11.
We will prove a suitable generalization of this theorem that holds for all weights and for all discriminants $D<0$ (Theorem 8.1).
Remark 1.12.
This theorem implies the supercongruences satisfied by $\frac{E_{4}}{j}$ and $\frac{E_{4}}{j-1728}$, namely Theorem 1.2, by choosing $k=4$ and D=3D=-3 or D=4D=-4 respectively. If one chooses $k=6$ and D=3D=-3, then one obtains the following nice supercongruence for $\frac{E_{6}}{j}$: for all primes $p\geq 5$ and all $n,l\in\mathbb{Z}^{+}$,
In this case, this theorem also implies that $\frac{384E_{6}}{j}$ is $2$-magnetic, though the $2$-magnetic property of $\frac{E_{6}}{j}$ was already proved in [PZ22, Theorem 2]. We refer to Example 5.4 for more examples of magnetic modular forms deduced from this theorem.
1.3. Structure of the paper
As mentioned before, the first four sections will only focus on the cases of weight $k\in\{4,6,8,10,14\}$, and the remaining four sections will concern general weights. We will also mostly focus on the case when the elliptic curve $C/\mathbb{Q}$ satisfies that $j(C)\notin\{0,1728\}$, and we will carry the assumption that $v_{p}(j(C))=0=v_{p}(j(C)-1728)$ when investigating the $p$-adic properties.
Section 2 will focus on the case of having a simple pole, where we raise the conjectural generalizations of Theorem 1.7 through a supersingular-ordinary dichotomy. We will also discuss the case when the simple pole is at a CM point, where one can already observe some three-term ASD congruences.
Section 3 will discuss the case of having a pole of higher order, and more specifically, will focus on the following space of modular forms
$\mathrm{MMF}_{k,C}\;:=\;\mathrm{Span}\bigg\{\frac{E_{k}}{j-j(C)},\;\frac{E_{k}}{(j-j(C))^{2}},\ldots,\frac{E_{k}}{(j-j(C))^{k-1}}\bigg\}.$
We will make comparisons between the space $\mathrm{MMF}_{k,C}$ with the $U_{p}$ action and the space $\mathrm{Sym}^{k-2}C$ with the Frobenius action, and as a generalization of 1.5, we will conjecture a suitable ASD congruence satisfied by modular forms in $\mathrm{MMF}_{k,C}$, which is given by the characteristic polynomial of the Frobenius action on $\mathrm{Sym}^{k-2}C$.
Section 4 will focus on the CM case. For a discriminant D<4D<-4 (of class number $1$), we will specifically focus on the case of having a pole (of order $\leq k-1$) at a CM point $\alpha_{D}$ of discriminant $D$, or equivalently, the space $\mathrm{MMF}_{k,D}:=\mathrm{MMF}_{k,C}$ when $C/\mathbb{Q}$ is a CM elliptic curve with CM by the imaginary quadratic order of discriminant $D$. We will construct a (rational) basis $\{G_{k,D}^{(r)}\}_{r=1}^{k-1}$ for the space $\mathrm{MMF}_{k,D}$, with each $G_{k,D}^{(r)}$ having a pole at $\alpha_{D}$ of order $r$, such that, roughly speaking, the basis is conjecturally
  1. (1)
    diagonal under the action of (powers of) $U_{p}$ for ordinary primes $p$, and
  2. (2)
    anti-diagonal under the action of (powers of) $U_{p}$ for supersingular primes $p$
via explicit supercongruences, where the (anti-)eigenvalues are described explicitly using the $p$-adic analogue of the Chowla–Selberg periods. Moreover, up to some nonzero integer multiple, each $G_{k,D}^{(r)}$ is conjecturally $(r^{\prime}-1)$-magnetic, where $r^{\prime}=\min(r,k-r)$. At the end of Section 4, we will also discuss the case when $D\in\{-3,-4\}$, or equivalently, $j(C)\in\{0,1728\}$.
The last four sections will focus on general weights and proofs of the main results. Section 5 will discuss the generalization to other weights via the help of relations and restate the main results and conjectures for general weights. Section 6 will recall the notations and preliminaries that are needed for the proofs of the main results. Section 7 will discuss the hypergeometric connection, namely Theorem 1.6, and use this to prove a generalization of Theorem 1.7 for all weights and for elliptic curves over number fields. Section 8 will discuss the proof of a generalization of Theorem 1.10 for all weights and all discriminants $D<0$ using the Borcherds–Shimura lift.
Throughout the paper, we also include many numerical examples under the help of LMFDB [LMF26] and PARI/GP [PAR23].
1.4. Related work
One of the main motivations of this paper is to obtain an explicit (though possibly conjectural) connection between meromorphic modular forms and symmetric powers of elliptic curves through numerically verifiable congruences. As mentioned before, the connection itself was already discussed in Brown–Fonseca [BF25], which built a comprehensive theory for meromorphic modular forms that is very closely related to this paper.
Indeed, in view of Kazalicki–Scholl [KS16], it is natural to expect the congruences that link meromorphic modular forms to symmetric powers of elliptic curves to be in the form of the ASD congruences. In this direction, the work of Allen–Long–Saad [ALS25] established certain ASD congruences for meromorphic modular forms on finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by extending the methods of Scholl [Sch85] and Kazalicki–Scholl [KS16]. They also treated the CM case in detail and deduced the existence of the (anti-)diagonal basis, though without explicitly identifying the basis and the (anti-)eigenvalues. Indeed, their main results should partially imply some of the congruences conjectured in this paper, at least when the weight $k\in\{4,6,8,10,14\}$ and when $p\geq k-1$. [^4] One should compare the appearances of $p\geq k-1$ here and below with the appearance of factorials in 5.3 and 5.8. However, it is not immediately clear whether the remaining cases follow as well, since our approach to general weights is different from theirs.
Another main issue that this paper encounters is the lack of Hecke operators on the space of meromorphic modular forms, or specifically, on the space $\mathrm{MMF}_{k,C}$. A reasonable workaround that one may expect is via the theory of $p$-adic modular forms. From this viewpoint, the work of Bordignon [Bor26] related meromorphic modular forms to overconvergent $p$-adic modular forms and described the explicit action of the $U_{p}$ operator via comparison theorems between rigid and de Rham cohomology, making it possible to deduce certain congruences of meromorphic modular forms by using the integrality arising from crystalline cohomology, and hence potentially resolving some of the conjectures in this paper when $p\geq k-1$.